Find the discrete time unit step response of the system with transfer function G(z) = (z−1)/z(z+1)
Answer #1


G(Z) = 2-1 z (2+1) 4(2) X(2) 2-1 2(2+1) z z Z-1 11 XL2) 1 1-2-1 x(0)2 urn) 2.T de X(2) (z Iransform) 4(2) 2 X(2) 2-1 Z(z+V z .2 +--7-2 د 2 ا+Z 1 7 + + 2 . اے 1 -2- Z Z -2-3 Z۔ 2. 3 -2- xh 2 + 2 - - - - ا = ( ع) 2 1 - 1 1 1 - 1 1 - 1 ۶ (۰) = 1 -